# ProRefine: Inference-Time Prompt Refinement with Textual Feedback

### v2

@@ Abstract
Agentic workflows, where multiple AI agents collaborate to accomplish complex tasks like reasoning or planning, play a substantial role in many cutting-edge commercial applications, and continue to fascinate researchers across nearly all fields for their potential to accomplish expensive, complex tasks that, until recently, only humans have been trusted to do. These workflows depend critically on the prompts used to provide the roles models play in such workflows. Poorly designed prompts that fail even slightly to guide individual agents can lead to sub-optimal performance that may snowball within a system of agents, limiting their reliability and scalability. To address this important problem of inference-time prompt optimization, we introduce ProRefine, an innovative inference-time optimization method that uses an agentic loop of LLMs to generate and apply textual feedback. ProRefine dynamically refines prompts for multi-step reasoning tasks without additional training or ground truth labels. Evaluated on five benchmark mathematical reasoning datasets, ProRefine significantly surpasses zero-shot Chain-of-Thought baselines by 3 to 37 percentage points. This approach not only boosts accuracy but also allows smaller models to approach the performance of their larger counterparts. This highlights its potential for building more cost-effective and powerful hybrid AI systems, thereby democratizing access to high-performing AI.
Introduction
The rapid evolution of Large Language Models (LLMs) has transformed applications from search engines to automated reasoning systems (Feng et al. 2024). Yet, their efficacy criticality depends on alignment with human values and preferences, a process often bottlenecked by the scarcity and subjectivity of human feedback (Gray and Suri 2019; Kahneman, Sibony, and Sunstein 2021; Prabhakar, Griffiths, and McCoy 2024; Weerasooriya et al. 2023). LLMs, trained on vast datasets, utilize alignment techniques to generate more human-like and accurate responses (Kirk et al. 2024; Rao et al. 2023; Sorensen et al. 2024).
@@ Introduction
Recent advances have shifted toward leveraging LLMs themselves as judges and evaluators (Bavaresco et al. 2024; Chiang and Lee 2023; Li et al. 2024; Liu et al. 2023; Wang et al. 2023; Zheng et al. 2023). Building on this, compound AI systems that use LLMs to generate feedback / critiques (Pryzant et al. 2023; Saunders et al. 2022). Akyurek et al. (2023) and Wadhwa et al. (2024) explore agentic frameworks to detect, critique, and refine tasks that require factual accuracy. Another approach is to iteratively refine the output using self-generated feedback (Madaan et al. 2023). TextGrad (Yuksekgonul et al. 2024) introduced automatic “differentiation” via text, using textual feedback to optimize the performance of individual components within a compound AI system.
Our work focuses on optimizing the prompt, a key element in chain-of-thought (CoT) (Wei et al. 2022) based LLM reasoning. CoT mimics human problem-solving by breaking down complex tasks into smaller, manageable steps. The system we introduce here, ProRefine (Inference-time Prompt Refinement with Textual Feedback), builds upon CoT by adaptively improving prompts using feedback ($LLM_{feedback}$) and an optimizer ($LLM_{optimizer}$) to refine prompts for the task-performing LLM ($LLM_{task}$). This workflow (Figure 1), motivated by the teacher-student framework (Torrey and Taylor 2013) where a teacher agent guides a student agent to perform a task by providing feedback at intermediate steps, but implemented via LLM interactions without pre-training, represents a novel approach to adaptive agentic reasoning. We explore policy optimization for aligning compound AI systems, drawing inspiration from TextGrad and policy gradient algorithms like PPO. This work investigates the following research questions.
RQ1
How effectively can textual feedback enhance LLM performance of LLMs during inference?
RQ2
RQ3
Figure 2: ProRefine example. Given an input query (which in this case has correct answer: 13), and an initial prompt, the task model ($LLM_{task}$) gives an incorrect answer. ProRefine uses two additional models, $LLM_{feedback}$ and $LLM_{optimizer}$, to iteratively improve the prompt as the $LLM_{task}$ generates its response. Refining the prompt during generation allows the feedback model to target local regions of the response, providing finer-grained feedback than waiting for the response to complete. We also provide an additional example illustrating our approach in Figure 5.
Although prior work has explored prompt optimization (Deng et al. 2022; Dong et al. 2024; Shin et al. 2020; Yang et al. 2024), they all often focus on either offline fine-tuning, which requires extensive training data, or universal application of largest, most capable models to every task. This presents a practical dilemma in many real-world scenarios. Continuously fine-tuning is not always feasible, and relying exclusively on state-of-the-art models is often computationally prohibitive. A different approach is needed for scenarios that require dynamic, on-the-fly repair for specific and difficult queries where a standard prompt fails. This is particularly true in resource-aware deployments, where a smaller model may suffice for most tasks but requires enhancement for a small subset of critical queries. The goal, therefore, shifts from finding a single, universally optimal prompt to performing targeted, inference-time intervention.
@@
The ability to break complex tasks into smaller steps and dynamically improve prompts offers a crucial advantage in multi-step agentic workflows where errors can compound. As illustrated in Figure 2, this method is also suitable for black-box LLMs where only API access is available. ProRefine could prove to be crucial in situations demanding greater interpretability, where feedback steps (outputs of $LLM_{feedback}$) offer insights into the reasoning correction process and applications requiring dynamic adaptation without retraining/fine-tuning cycles. To demonstrate its effectiveness, we evaluate ProRefine across five benchmark mathematical reasoning datasets, showing it offers a robust alternative to solely scaling up the base model for all queries.
@@
  • We propose a novel method - ProRefine - for prompt optimization inference-time using textual feedback.
@@
Essentially, ProRefine is an elegant, training-free, inference-time solution that fixes the core prompt to improve reasoning, making it suitable for practical hybrid-model deployment.
Related Work
ProRefine draws inspiration from and contributes to several interconnected research areas. The performance of LLMs is heavily dependent on the quality of the prompts they receive. Early efforts in this domain centered on manual prompt crafting (Wei et al. 2022), a meticulous process of designing effective prompts to elicit desired responses. Recognizing the limitations and scalability challenges of manual methods, research has increasingly focused on automatic prompt optimization with a growing emphasis on agentic workflows that enable dynamic and adaptive reasoning.
Prompt Generation
Some pioneering automatic methods, such as AutoPrompt (Shin et al. 2020) and RLPrompt (Deng et al. 2022), employ gradient-based search and reinforcement learning techniques, respectively. AutoPrompt (Shin et al. 2020) uses gradient-based search to generate prompts for masked language models. It reformulates tasks as fill-in-the-blank problems, achieving performance comparable to supervised models in tasks like sentiment analysis. However, it requires training data and gradient access, limiting its applicability to black-box models. Other approaches leverage LLMs themselves for prompt generation (Mehta, Pyatkin, and Srikumar 2024; Pryzant et al. 2023; Yang et al. 2024, 2022; Zhou et al. 2022). Recent works like Promptomatix (Murthy et al. 2025) and EvoAgentX (Wang et al. 2025) extend this direction by enabling automatic prompt refinement across multiple tasks, workflows, and tools. ProRefine distinguishes itself by operating simply at inference-time, requiring no training data, gradient access, or model retraining, while enabling prompt refinement in dynamically evolving settings.
Self-Refinement
There is a substantial and growing body of work exploring the capacity of LLMs to act as judges or evaluators (Bavaresco et al. 2024; Chiang and Lee 2023; Li et al. 2024; Liu et al. 2023; Verga et al. 2024; Wang et al. 2023; Zheng et al. 2023; Zhuge et al. 2024). This capability has been leveraged to assess response quality or provide self-feedback. ProRefine adopts this principle, using LLM-generated textual feedback to improve its own prompting process. Unlike prior uses of LLM evaluation solely for ranking or filtering, ProRefine uses that feedback in a closed-loop for optimization during inference.
The idea of LLM iterative refinement is highly relevant. Self-Refine (Madaan et al. 2023) is a prominent example, where an LLM generates both output and feedback, using the latter for refinement. ARIES (Zeng et al. 2025) further enhances refinement via Elo-style agent debate. Other works explore self-critiquing (Saunders et al. 2022) and reinforcement learning for critique generation (RL4F) (Akyurek et al. 2023), along with various feedback and refinement mechanisms (Dong et al. 2024; Khattab et al. 2024; Qu et al. 2024; Ranaldi and Freitas 2024; Schick et al. 2023; Wadhwa et al. 2024), and Monte Carlo-based refinement in math reasoning (MC-NEST) (Rabby, Keya, and Auer 2025). While ProRefine shares the self-refinement spirit, it focuses on prompt refinement, suitable for agentic workflows and black-box LLMs, while avoiding reinforcement learning and direct output modification.
Inference-Time Scaling
ProRefine belongs to the broader category of inference-time methods (Muennighoff et al. 2025; Snell et al. 2024a), that improve LLMs without without weight modification (Du et al. 2024). Inference-time methods aim to improve the performance of models by utilizing test-time compute resources. TextGrad (Yuksekgonul et al. 2024) performs gradient-free inference-time optimization using textual feedback. ProRefine applies a similar idea to intermediate prompt refinement for dynamic reasoning chains. TextGrad relies on supervised fine-tuning, whereas ProRefine operates without training data, offering ease of integration. Other inference-time strategies include RL-of-Thoughts (Hao et al. 2025) and Reward-Is-Enough (Song et al. 2025), which apply RL-based signal propagation during inference. AvR (Alignment via Refinement) (Zhang et al. 2025a) proposes recursive CoT refinement using long-form reasoning. ProRefine, by contrast, performs step-level feedback on prompts rather than final outputs, and requires no external tools or supervision.
Agentic Workflows
ProRefine also fits into a broader trend toward agentic workflows. AFlow (Zhang et al. 2025b) automates agentic workflows through prompt-based search over prior structures, while EvoAgentX (Wang et al. 2025) evolves agent behaviors and topologies. Meanwhile, Mass (Zhou et al. 2025) and DebFlow (Su et al. 2025) optimize multi-agent configurations via interleaved search and debate. ProRefine focuses instead on optimizing individual agent prompts within fixed workflows, complementing these methods. Unlike tool-integrated or debate-based systems, ProRefine remains model-agnostic and easy to integrate into any prompt-based agent loop.
@@
ProRefine is an inference-time prompt optimization algorithm that optimizes prompts by using textual feedback. ProRefine involves interactions between three LLMs:
@@
$LLM_{feedback}$: A model that critiques the $LLM_{task}$’s output, providing detailed feedback on improvements. This model should be capable of providing insightful and accurate critiques (Bai et al. 2022; Saunders et al. 2022).
@@
(Algorithm 1) works as follows:
@@
 if EOS_tokenEOS\_token in $o_{i}$ then
    break
@@
Unifying Verifier and Feedback
At inference time, verifiers play a crucial role in judging model outputs (Cobbe et al. 2021; Lightman et al. 2024; Snell et al. 2024b). Often trained on extensive corpora, these verifiers can be outcome-based, assessing the final result’s correctness, or process-based, evaluating the validity of intermediate steps (Lightman et al. 2024; Snell et al. 2024b; Uesato et al. 2022). For simplicity in this study, we do not train a bespoke verifier; rather, we employ the Llama3.1-70B-instruct model to function as both the feedback mechanism ($LLM_{feedback}$) and the verifier. We manage these roles through separate API calls, each with a role-defining prompt. A smaller model, specifically fine-tuned for these tasks, could also be used. The verifier’s function is to evaluate the initial output generated by $LLM_{task}$ for each query. If verifier assess this output is incorrect, the refinement process is triggered; otherwise, the output is used as is. This also saves computation on answers that are already correct.
@@
Implementation Details
The granularity and duration of the optimization process are governed by two parameters: $k$, the number of tokens per step, and $n$, the maximum number of steps. These parameters can be adjusted based on the complexity of the task and the desired output quality. For example, rather than generating feedback every $k$ tokens, we might instead choose to provide feedback after each sentence or paragraph, particularly in tasks such as machine translation or text summarization, where larger semantic units may be more meaningful. We run all of our experiments on compute nodes with a single Nvidia A100 GPU (80GB), 24 core processor, and 220GB RAM. For model’s generate function we set the following parameters: num_return_sequences=1, do_sample=False, top_p=None, temperature=None.
Experiments and Evaluation
Data
We evaluate ProRefine on five reasoning tasks, each of which involves multi-step reasoning, making them suitable for evaluating prompt optimization in agentic workflows. We include the original dataset split sizes in (train/validation/test) format: object counting and word sorting from the BIG-Bench Hard benchmark (Srivastava et al. 2023) (50/100/100), grade-school math problem-solving from GSM8K (Cobbe et al. 2021) (200/300/1319), math word problems from SVAMP (Patel, Bhattamishra, and Goyal 2021) (2516/622/1000), and algebraic word problems from AQUARAT (Ling et al. 2017) (97467/254/254). We use the same splits and evaluation as Yuksekgonul et al. (2024) for object counting, word sorting, and GSM8K.
Experimental Setup

### v3

@@ Abstract
Agentic workflows, where multiple AI agents collaborate to accomplish complex tasks like reasoning or planning, play a substantial role in many cutting-edge commercial applications, and continue to fascinate researchers across fields for their potential to accomplish expensive, complex tasks that, until recently, only humans have been trusted to do. These workflows depend critically on the prompts used to provide the roles models play in such workflows. Poorly designed prompts that fail even slightly to guide individual agents can lead to sub-optimal performance that may snowball within a system of agents, limiting their reliability and scalability. To address this important problem of inference-time prompt optimization, we introduce ProRefine, an innovative inference-time optimization method that uses an agentic loop of LLMs to generate and apply textual feedback. ProRefine dynamically refines prompts for multi-step reasoning tasks without additional training or ground truth labels. Evaluated on five benchmark mathematical reasoning datasets, ProRefine significantly surpasses zero-shot Chain-of-Thought baselines by 3 to 37 percentage points. This approach not only boosts accuracy but also allows smaller models to approach the performance of their larger counterparts. This highlights its potential for building more cost-effective and powerful hybrid AI systems, thereby democratizing access to high-performing AI.
1 Introduction
The advancement of Large Language Models (LLMs) is intrinsically linked to their alignment with human values and preferences (Feng et al., 2024). While Reinforcement Learning from Human Feedback (RLHF) has been the cornerstone of this effort (Christiano et al., 2017), recent research has pivoted towards using LLMs themselves as scalable proxies for human judgment, serving as evaluators, critics, and sources of feedback Zheng et al. (2023); Pryzant et al. (2023); Saunders et al. (2022). This has given rise to sophisticated agentic frameworks that can detect errors, critique outputs, and iteratively refine them, particularly for tasks demanding factual correctness Akyurek et al. (2023); Madaan et al. (2023). Methods like TextGrad have even demonstrated how textual feedback can “differentiate” through complex systems to optimize performance Yuksekgonul et al. (2024).
@@ 1 Introduction
Our work focuses on optimizing the prompt, a key element in chain-of-thought (CoT) (Wei et al., 2022) based LLM reasoning. Although prior work has explored prompt optimization Deng et al. (2022); Dong et al. (2024); Shin et al. (2020); Yang et al. (2024), they all often focus on either offline fine-tuning, which requires extensive training data, or universal application of largest, most capable models to every task. This presents a practical dilemma in many real-world scenarios. Continuously fine-tuning is not always feasible, and relying exclusively on state-of-the-art models is often computationally prohibitive. A different approach is needed for scenarios that require dynamic, on-the-fly repair for specific and difficult queries where a standard prompt fails. This is particularly true in resource-aware deployments, where a smaller model may suffice for most tasks but requires enhancement for a small subset of critical queries. The goal, therefore, shifts from finding a single, universally optimal prompt to performing targeted, inference-time intervention.
To address this need, we introduce ProRefine (Inference-time Prompt Refinement with Textual Feedback), which builds upon CoT by adaptively improving prompts using feedback ($LLM_{feedback}$) and an optimizer ($LLM_{optimizer}$) to refine prompts for the task-performing LLM ($LLM_{task}$). This workflow (Figure 1), motivated by the teacher-student framework Torrey and Taylor (2013) where a teacher agent guides a student agent to perform a task by providing feedback at intermediate steps, but implemented via LLM interactions without pre-training, represents a novel approach to adaptive agentic reasoning. We explore policy optimization for aligning compound AI systems, drawing inspiration from TextGrad and policy gradient algorithms, such as PPO.
@@
The ability to break complex tasks into smaller steps and dynamically improve prompts offers a crucial advantage in multi-step agentic workflows where errors can compound. As illustrated in Figure 4 in the Appendix, This method is also suitable for black-box LLMs where only API access is available. ProRefine could prove to be crucial in situations demanding greater interpretability, where feedback steps (outputs of $LLM_{feedback}$) offer insights into the reasoning correction process and applications requiring dynamic adaptation without retraining/fine-tuning cycles. To demonstrate its effectiveness, we evaluate ProRefine across five benchmark mathematical reasoning datasets, showing it offers a robust alternative to solely scaling up the base model for all queries.
@@
  • We propose a novel method - ProRefine - for prompt optimization at inference-time using textual feedback.
@@
2 ProRefine
@@
is an inference-time prompt optimization algorithm that optimizes prompts by using textual feedback. ProRefine involves interactions between three LLMs:
@@
$LLM_{feedback}$: A model that critiques the $LLM_{task}$’s output, providing detailed feedback on improvements. This model should be capable of providing insightful and accurate critiques Bai et al. (2022); Saunders et al. (2022).
@@
(Algorithm 1) works as follows [^1] The code is available at https://github.com/deepakpandita57/ProRefine_public:
@@ 1 $p^{*}=p$
3 if EOS_tokenEOS\_token in $o_{i}$ then
4    break
@@
The granularity and duration of the optimization process are governed by two parameters: $k$, the number of tokens per step, and $n$, the maximum number of steps. These parameters can be adjusted according to the task’s complexity and the desired output quality. For example, rather than generating feedback every $k$ tokens, we might instead choose to provide feedback after each sentence or paragraph, particularly in tasks such as machine translation or text summarization, where larger semantic units may be more meaningful.
Unifying Verifier and Feedback:
At inference time, verifiers play a crucial role in judging model outputs Cobbe et al. (2021); Lightman et al. (2024); Snell et al. (2024b). For simplicity in this study, we do not train a bespoke verifier; rather, we employ the Llama3.1-70B-instruct model to function as both the feedback mechanism ($LLM_{feedback}$) and the verifier. We manage these roles through separate API calls, each with a role-defining prompt. A smaller model, specifically fine-tuned for these tasks, could also be used. The verifier’s function is to evaluate the initial output generated by $LLM_{task}$ for each query. If the verifier assesses the output to be incorrect, the refinement process is triggered; otherwise, the output is used as is. This also saves computation on answers that are already correct.
@@
3 Experiments and Evaluation
@@
3.1 Data
We evaluate ProRefine on five reasoning tasks, each of which involves multi-step reasoning, making them suitable for evaluating prompt optimization in agentic workflows. We utilize object counting and word sorting from the BIG-Bench Hard benchmark Srivastava et al. (2023), grade-school math problem-solving from GSM8K Cobbe et al. (2021), math word problems from SVAMP Patel et al. (2021), and algebraic word problems from AQUARAT Ling et al. (2017). See Appendix 5 for details about data splits.
3.2 Experimental Setup
We experiment with three models - Llama3.2-1B-instruct, Llama3.2-3B-instruct, and Llama3.1-8B-instruct Meta (2024) for $LLM_{task}$. The prompts are optimized using ProRefine, with Llama3.1-70B-instruct used for feedback generation, prompt optimization, and evaluation. We select the values of hyperparameters $k=10$ and $n=25$ to control the granularity of feedback and duration of optimization. Hyperparameters $k$ and $n$ were fixed based on general preliminary exploration and not tuned per task using benchmark training/validation data.
We compare ProRefine against the zero-shot Chain-of-Thought (CoT) baseline and TextGrad Yuksekgonul et al. (2024), and report test accuracy with 95% confidence interval. It is essential to remember that TextGrad is a supervised fine-tuning method that utilizes both the training and validation sets.
3.3 Results
Object Counting:
Word Sorting:
GSM8K:
SVAMP:
AQUARAT:
Gains over CoT range from 0.920.9-2 percentage points, but declines of 5.4225.4-22 points are also observed. ProRefine exceeds TextGrad on 2 of 3 models, with 2.74.72.7-4.7 percentage point gains, although performance drops of 10.618.110.6-18.1 points are also recorded.
Our results demonstrate that using ProRefine with an optimal verifier significantly improves performance for all tasks, achieving the best results in 13 out of 15 cases, highlighting the critical role of verifier quality. Notably, the number of significant improvements increases with larger model sizes. We also observe that ProRefine enables smaller models, such as Llama3.2-3B-instruct and Llama3.1-8B-instruct, to approach the zero-shot performance of larger models like Llama3.1-8B-instruct and Llama3.1-70B-instruct, respectively.
4 Related Work
ProRefine draws inspiration from and contributes to several interconnected research areas. The performance of LLMs is heavily dependent on the quality of the prompts they receive. Early efforts in this domain centered on crafting prompts manually Wei et al. (2022), a meticulous process of designing effective prompts to elicit desired responses. Recognizing the limitations and scalability challenges of manual methods, research has increasingly focused on automatic prompt optimization with a growing emphasis on agentic workflows that enable dynamic and adaptive reasoning.
Prompt Generation:
Some pioneering automatic methods, such as AutoPrompt Shin et al. (2020) and RLPrompt Deng et al. (2022), employ gradient-based search and reinforcement learning techniques, respectively. AutoPrompt Shin et al. (2020) uses gradient-based search to generate prompts for masked language models. It reformulates tasks as fill-in-the-blank problems, achieving performance comparable to supervised models in tasks like sentiment analysis. However, it requires training data and gradient access, limiting its applicability to black-box models. Other approaches leverage LLMs themselves for prompt generation Mehta et al. (2024); Pryzant et al. (2023); Yang et al. (2024, 2022); Zhou et al. (2022). Recent works like Promptomatix Murthy et al. (2025) and EvoAgentX Wang et al. (2025) extend this direction by enabling automatic prompt refinement across multiple tasks, workflows, and tools. ProRefine distinguishes itself by operating simply at inference-time, requiring no training data, gradient access, or model retraining, while enabling prompt refinement in dynamically evolving settings.
Self-Refinement:
There is a substantial and growing body of work exploring the capacity of LLMs to act as judges or evaluators Bavaresco et al. (2024); Chiang and Lee (2023); Li et al. (2024); Liu et al. (2023); Verga et al. (2024); Wang et al. (2023); Zheng et al. (2023); Zhuge et al. (2024). This capability has been leveraged to assess response quality or provide self-feedback. ProRefine adopts this principle, using LLM-generated textual feedback to improve its own prompting process. Unlike prior uses of LLM evaluation solely for ranking or filtering, ProRefine uses that feedback in a closed-loop for optimization during inference.
The idea of LLM iterative refinement is highly relevant. Self-Refine Madaan et al. (2023) is a prominent example, where an LLM generates both output and feedback, using the latter for refinement. ARIES Zeng et al. (2025) further enhances refinement via Elo-style agent debate. Other works explore self-critiquing Saunders et al. (2022) and reinforcement learning for critique generation (RL4F) Akyurek et al. (2023), along with various feedback and refinement mechanisms Dong et al. (2024); Khattab et al. (2024); Qu et al. (2024); Ranaldi and Freitas (2024); Schick et al. (2023); Wadhwa et al. (2024), and Monte Carlo-based refinement in math reasoning (MC-NEST) Rabby et al. (2025). While ProRefine shares the self-refinement spirit, it focuses on prompt refinement, suitable for agentic workflows and black-box LLMs, while avoiding reinforcement learning and direct output modification.
Inference-Time Scaling:
ProRefine belongs to the broader category of inference-time methods Muennighoff et al. (2025); Snell et al. (2024a), that improve LLMs without without weight modification Du et al. (2024). Inference-time methods aim to improve the performance of models by utilizing test-time compute resources. TextGrad Yuksekgonul et al. (2024) performs gradient-free inference-time optimization using textual feedback. ProRefine applies a similar idea to intermediate prompt refinement for dynamic reasoning chains. TextGrad relies on supervised fine-tuning, whereas ProRefine operates without training data, offering ease of integration. Other inference-time strategies include RL-of-Thoughts Hao et al. (2025) and Reward-Is-Enough Song et al. (2025), which apply RL-based signal propagation during inference. AvR (Alignment via Refinement) Zhang et al. (2025a) proposes recursive CoT refinement using long-form reasoning. ProRefine, by contrast, performs step-level feedback on prompts rather than final outputs, and requires no external tools or supervision.
Agentic Workflows:
ProRefine also fits into a broader trend toward agentic workflows. AFlow Zhang et al. (2025b) automates agentic workflows through prompt-based search over prior structures, while EvoAgentX Wang et al. (2025) evolves agent behaviors and topologies. Meanwhile, Mass Zhou et al. (2025) and DebFlow Su et al. (2025) optimize multi-agent configurations via interleaved search and debate. ProRefine focuses instead on optimizing individual agent prompts within fixed workflows, complementing these methods. Unlike tool-integrated or debate-based systems, ProRefine remains model-agnostic and easy to integrate into any prompt-based agent loop.
5 Experiments and Evaluation
5.1 Data
We evaluate ProRefine on five reasoning tasks, each of which involves multi-step reasoning, making them suitable for evaluating prompt optimization in agentic workflows. We include the original dataset split sizes in (train/validation/test) format: object counting and word sorting from the BIG-Bench Hard benchmark Srivastava et al. (2023) (50/100/100), grade-school math problem-solving from GSM8K Cobbe et al. (2021) (200/300/1319), math word problems from SVAMP Patel et al. (2021) (2516/622/1000), and algebraic word problems from AQUARAT Ling et al. (2017) (97467/254/254). We use the same splits and evaluation as Yuksekgonul et al. (2024) for object counting, word sorting, and GSM8K.
5.2 Experimental Setup